OpenStudy (mony01):

Simplify the expression? tan(sin^-1(x))

OpenStudy (anonymous):

first, change tan and sin^-1 \[\tan =\frac{ \sin }{ \cos }\] and sin^-1 = \[\frac{ 1 }{ \sin }\]

OpenStudy (anonymous):

so, you can rewrite them like this: \[\frac{ \sin }{ \cos } \left( \frac{ 1 }{ \sin } (x)\right)\]

OpenStudy (anonymous):

sin^-1x is arcsin(x), it doesn't equal 1/sinx.

OpenStudy (anonymous):

i hope this will help you..

OpenStudy (anonymous):